How much work is done by the gravitational force when a 265-kg pile driver falls 2.80 m?
1
Expert's answer
2013-11-22T09:21:55-0500
lets say g=10m/s^2 (10N/kg)
then
A=F*s=m*g*s=265kg*10N/kg*2.8m=7420 J

A=7420 J